OpenStudy (anonymous):

2x+2x+6x+24=0 whats x?

OpenStudy (anonymous):

so group like terms, all the x's add up to 10x and isolate that by subtracting 24 from both sides. Divide everything by 10 and you get x=-2.4.

jhonyy9 (jhonyy9):

10x+24=0 10x=-24 x=-24/10 x=-12/5